40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 7 14:39:49 2013, Jenne, Update, LSC, Lock Acquisition Game Plan LockAcquisition_7Nov2013.pdf
    Reply  Fri Nov 8 18:12:21 2013, Jenne, Update, LSC, PRFPMI: Not crossing any resonances 2f_resonances_Xarm_CARMsweep.pngAllModFreq_resonances_Xarm_CARMsweep.png
    Reply  Tue Nov 12 16:49:22 2013, Jenne, Update, LSC, Xend QPD and Whitening board pulled IMG_1262.JPGIMG_1260.JPGIMG_1264.JPGIMG_1276.JPG
       Reply  Wed Nov 20 17:05:15 2013, Jenne, Update, LSC, Xend QPD and Whitening board replaced 
       Reply  Wed Dec 18 00:26:15 2013, Jenne, Update, LSC, Xend QPD schematic investigation 
          Reply  Wed Dec 18 11:32:34 2013, rana, Update, LSC, Xend QPD schematic investigation 
    Reply  Wed Nov 13 01:35:40 2013, Jenne, Update, LSC, PRM motion causing trouble? xarmRIN_vs_CARM.pngASportOverlap_moveGaussianBeams.png
       Reply  Wed Nov 13 09:31:15 2013, Gabriele, Update, LSC, PRM motion causing trouble? 
          Reply  Wed Nov 13 19:22:58 2013, Jenne, Update, LSC, PRM motion correlated to intracavity power POPDC_sensors_30oct2013.pngAOMmodulated_POPDC_OSEM.pdf
             Reply  Thu Nov 14 02:50:43 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF PRM_OL_TF_model.pngPRM_OLPIT_TF.pdfPRM_OLYAW_TF.pdf
                Reply  Mon Nov 18 21:02:54 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF PRM_OLG_NewOld.pngPRM_oplev_improvement.pdf
                   Reply  Mon Nov 18 21:40:24 2013, Koji, Update, LSC, PRM oplev measured and modeled TF 
                      Reply  Tue Nov 19 14:47:44 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF 
             Reply  Thu Nov 14 02:55:26 2013, rana, Update, SUS, PRM motion correlated to intracavity power 
                Reply  Thu Nov 14 11:41:19 2013, Steve, Update, SUS, PRM sensors effected by IR IRon_off.png
                   Reply  Thu Nov 14 14:35:12 2013, Steve, Update, SUS, IR effect on MC and PRM sensors MC_PRM_IReffect.png
                      Reply  Fri Dec 6 10:03:07 2013, Steve, Update, SUS, IR effect on MC sensors only 10minPMCnotLocked.png6Dec2013.png
Message ID: 9367     Entry time: Tue Nov 12 16:49:22 2013     In reply to: 9360     Reply to this: 9418   9484
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Xend QPD and Whitening board pulled 

Quote:

* Whitening for the transmission QPDs needs to be thought about more carefully. (Calculation, then hardware)

 I have the X end transmission QPD, as well as the whitening board, out on the electronics bench.  Since the Thorlabs high-gain TRX PD also goes through this whitening board, we have no transmission signal for the Xarm at this time. The whitening board was in the left-most slot, of the top crate in the Xend rack.  The only cables that exist for it (like the Yend), are the ribbon from the QPD, the 4-pin lemo from the Thorlabs PD, and the ribbon going to the ADC.

I have taken photos, and want to make sure that I know what is going on on the circuits, before I put them back in. 

The QPD:

IMG_1262.JPG

IMG_1260.JPG

The whitening board:

IMG_1264.JPG

IMG_1276.JPG

ELOG V3.1.3-